OpenStudy (superhelp101):

|x + 3| + 7 < 10 describe the graph in a complete sentence

OpenStudy (superhelp101):

It should actually be \[\le \]

OpenStudy (superhelp101):

any help @abb0t @Ashleyisakitty @bishoy41 @tyara45 @This_Is_Batman @lraquel @qdbug @radar @robtobey any help plzzzzz :(

OpenStudy (anonymous):

well its just adding and subtracting

OpenStudy (superhelp101):

I just don't get the graph tho

OpenStudy (anonymous):

well you have the equation |x + 3| + 7 <= 10 you first have to look for x. l x + 3 l <= 3 and because this is a absolute value equation you must make it into 2 equations x + 3 <= 3 x + 3 >= -3 after that you just grapgh what you have

OpenStudy (superhelp101):

yes I just have problems with graphing

OpenStudy (vishweshshrimali5):

Perfect answer @FriedRice . Great work

OpenStudy (vishweshshrimali5):

Lets see the graph part, then.

OpenStudy (superhelp101):

x≥-6

OpenStudy (superhelp101):

x≤0

OpenStudy (vishweshshrimali5):

As @FriedRice mentioned, you have two equations: (1) x+3 <= 3 => x <= 0 (2) x+3 >= -3 => x>= -6 Correct @superhelp101 :)

OpenStudy (superhelp101):

I can't explain the graph for them

OpenStudy (vishweshshrimali5):

Now for drawing the feasible region (the region represented by inequalities): (1) Change the inequality sign into = sign and then draw the graph.

OpenStudy (superhelp101):

okay

OpenStudy (vishweshshrimali5):

I am going to draw the graph of 1 inequality and then you will draw the other one. So, which one should I help you with first ?

OpenStudy (superhelp101):

I think I get it know. I had problems with shading but I understand! Thank You @vishweshshrimali5

OpenStudy (superhelp101):

And @FriedRice

OpenStudy (vishweshshrimali5):

No problem. It was pleasure to help you out.